# Can ChatGPT Detect DeepFakes? A Study of Using Multimodal Large Language Models for Media Forensics

### v1

@@ 1 Introduction
The meteoric rise of Generative AI (GenAI) models is one of the most exciting developments in recent years. State-of-the-art GenAI models have demonstrated incredible abilities to create realistic images, audio, and videos111e.g., Midjourney, www.midjourney.com, Stable Diffusion https://stability.ai/ for image generation, Elevenlab, https://elevenlabs.io/ for audio generation, and Pika, https://pika.art/, Txt2vid, https://github.com/tpulkit/txt2vid, and OpenAI’s Sora, https://openai.com/sora for video generation. from text prompts. While AI-generated content has numerous beneficial uses, such as in the movie and advertising industry, its misuse to produce deleterious content, commonly known as DeepFakes, seriously undermines the credibility of information and trust in digital media. As a result, identifying DeepFakes has become a crucial and timely task in media forensics.
@@
We hope that this study will encourage future exploration of the use and improvement of LLMs for media forensics and DeepFake detection.
The remainder of the paper is organized as follows. Section 2 provides an overview of the relevant literature on LLMs and deep-fake face detection. Section 3 presents the methodology of our study, and an in-depth evaluation and analysis of our main experimental results are given in Section 4. Section 5 concludes the article with proposals of potential strategies to enhance LLM performance in media forensic tasks.
@@ 2.2 DeepFake Faces: Generation and Detection
Refer to captionRefer to captionRefer to captionRefer to captionRefer to caption
Refer to captionRefer to captionRefer to captionRefer to captionRefer to caption
@@
Figure 2: Which of these images are real and AI-generated?
Answer:
@@ 3 Methodology
Data: Our experiments are based on a set of $1,000$ real face images from the FFHQ dataset [18] dataset and another $2,000$ images created with generative AI models from the $DF^{3}$ dataset [16]. All images contain a single human face. Two AI generative models are considered, namely StyleGAN2 [19] and Latent Diffusion [28]. We also adopt two evaluation protocols from the $DF^{3}$ dataset [16]. This includes assessing the basic detection performance of raw data and evaluating the robustness of post-processed data through mixed operations such as JPEG Compression, Gaussian Blur, face blending, adversarial attacks, and multi-image compression. Detailed information on the data used is given in Table 1. A few examples of the real and AI-generated images are shown in Fig. 3.
   Real   (FFHQ)
SG2 (Raw)
LD(Raw)
   SG2  (PP’ed)
   LD   (PP’ed)
Refer to captionRefer to captionRefer to captionRefer to captionRefer to caption
Refer to captionRefer to captionRefer to captionRefer to captionRefer to caption
Refer to captionRefer to captionRefer to captionRefer to captionRefer to caption
@@
Table 1: Detailed information of the evaluation dataset from $DF^{3}$ [16].
Real (FFHQ) Raw data Post-processed
StyleGAN2 Latent Diffusion StyleGAN2 Latent Diffusion
Number 1,000 1,000 1,000 1,000 1,000
Image Size 512$\times$512 512$\times$512 512$\times$512 256$\times$256 256$\times$256
Format PNG PNG JPEG PNG, JPEG PNG, JPEG
@@
Figure 4: Examples of GPT4V for DeepFake face detection. Left: Results for AI-generated images from the $DF^{3}$ dataset [16]. Right: Results for real faces from the FFHQ dataset [18]. The responses for AI-generated faces are labeled in pink, while those for the real faces are labeled in green. Both success (w/ Refer to caption) and failure (w/ Refer to caption) cases are shown. See texts for details.
@@
Performance Metric: For each text-image prompt, we query the LLM multiple times and calculate a numerical decision score as the average of the results (No $=0$ and Yes $=1$). This approach offers two benefits. Firstly, it diminishes the variability in LLM responses to identical queries, attributable to the probabilistic nature of the underlying LLMs. Secondly, using numerical decision scores enables the application of performance metrics beyond mere accuracy, such as the classification accuracies and the area under the ROC (AUC) score. Compared to classification accuracies, the AUC score is less affected by imbalanced data, provides a more comprehensive performance evaluation, and allows us to compare the LLM’s performance with existing programmed detection methods. AUC score is a real number in $[0,1]$, with higher values corresponding to better performance. As the LLM may decline to respond to a query, another important performance metric is the rejection rate, which measures the fraction of queries that the LLM declines.
@@ 4.1 Qualitative and Quantitative Results
The quantitative results corroborate this observation. In Fig. 6, we show the receiver operational curves (ROCs) and the corresponding AUC scores obtained using API calls (as described in Section 3) over the evaluation dataset with the same prompt – GPT4V has a $79.5\%$ AUC score on raw latent diffusion-generated face images and $77.2\%$ AUC score on StyleGAN-generated face images. The performance confirms that the GPT4V model obviously did not make random guesses on this task (corresponding to a ROC as a diagonal line and a $50\%$ AUC score). Compared to the GPT4V model, the Gemini model shows a slight decrease in performance.
@@
Figure 6: ROC curves of GPT4V and Gemini 1.0 Pro on the DeepFake detection based on averaging the predictions of five rounds of queries, (a) on raw data, (b) on post-processed data.
@@
On the other hand, we also note that most errors of GPT4V occur on real images – per Table 3, the classification accuracies on real images are around $50\%$, drastically different from those of AI-generated images, which is above $90\%$. Some intuitions can be obtained when we examine the real face images for these error cases, as shown in Fig. 4. These cases include semantic features unusual for “typical” face images. For instance, different age group (baby in the first case) or unique hair color (second case) to style (third case). This suggests that the semantic abnormality identified by GPT4V may not be specific to DeepFake faces. This problem may be solved by refining the model. In contrast, the Gemini model achieves a classification accuracy of $83.3\%$ on real images, dropping to around $50\%$ on generated faces. The examples in Fig. 5 show that the Gemini model’s response lacks rationality in analyzing the synthesis artifacts.
Table 3: Comparison of Accuracy (%) in detecting DeepFake faces. ‘SG2’ stands for the StyleGAN2 model, and ‘LD’ represents the Latent Diffusion model.
@@ 4.3 Improvements
So far, we have only tested simple queries. It has been demonstrated that using better prompts constructed with chain-of-thought prompts [35], few-shot prompting [40], which provide step-to-step guides in an interactive conversation with the LLM can elicit more relevant responses. The API interface of GPT4V and Gemini 1.0 does not allow multiple rounds of dialog because no consistent context is stored across API calls. Therefore, such interactive guidance can only be used with the web interface through a manual interaction, i.e., they cannot be automated using API calls. We provide two exploratory approaches in Fig. 10: one employs decomposed local images, a method we refer to as decomposition-based prompting, and the other utilizes a few-shot prompting technique. By supplying decomposed parts of images, we direct the model’s attention towards finer local patterns and reveal subtle visual anomalies. Additionally, using few-shot learning strategies with brief synthesis instructions imparts crucial forensic knowledge to the model, enabling the model to classify two samples that were previously misidentified correctly. These initial results suggest that using more crafted prompts can improve performance. However, we will wait for the LLMs to enable consistent API calls.
@@ 5 Conclusion
In this study, we investigate the potential of leveraging multimodal LLMs for tasks related to media forensics. Our research will broaden the application of multimodal LLMs to include a wider array of media forms, particularly focusing on video analysis. Rather than simply applying image-based detection techniques to video frames, a more integrated approach would involve direct video content processing. Additionally, we aim to enhance the detection of text-image mis-contextualization, where images and text are misleadingly paired to spread false information. Future endeavors will also explore developing more sophisticated prompting strategies and integrating these models with conventional signal or data-driven detection techniques.
@@
Acknowledgement. Siwei Lyu is supported by National Natural Science Foundation under grant No.62076213 SaTC-2153112. Xin Li is supported by National Natural Science Foundation under grant CCSS-2348046 and SUNY-Albany start-up funds. Baoyuan Wu is supported by National Natural Science Foundation of China under grant No.62076213, Shenzhen Science and Technology Program under grants No.RCYX20210609103057050, and the Longgang District Key Laboratory of Intelligent Digital Economy Security.

### v3

@@ 1 Introduction
The meteoric rise of Generative AI (GenAI) models is one of the most exciting developments in recent years. State-of-the-art GenAI models have demonstrated incredible abilities to create realistic images, audio, and videos111e.g., Midjourney www.midjourney.com and Stable Diffusion https://stability.ai/ for image generation, Elevenlab https://elevenlabs.io/ for audio generation, and Pika https://pika.art/ and OpenAI’s Sora https://openai.com/sora for video generation. from text prompts. While AI-generated content has numerous beneficial uses, such as in the movie and advertising industry, its misuse to produce deleterious content, commonly known as DeepFakes, seriously undermines the credibility of information and trust in digital media. As a result, identifying DeepFakes has become a crucial and timely task in media forensics.
@@
We hope that this study will encourage future exploration of the use and improvement of LLMs for media forensics and DeepFake detection. The remainder of the paper is organized as follows. Section 2 provides an overview of the relevant literature on LLMs and Deepfake face detection. Section 3 presents the methodology of our study. Comprehensive evaluation results and analysis are given in Section 4, and Section 5 concludes the article..
@@ 3 Methodology
Table 1: Detailed information of the evaluation dataset from $DF^{3}$ [17]. ‘SG2’ stands for the StyleGAN2 model, ‘LD’ represents the Latent Diffusion model, and ‘PP’ed’ means post-processed data.
Real Raw PP’ed
SG2 LD SG2 LD
Number 1,000 1,000 1,000 1,000 1,000
Image Size $512^{2}$ $512^{2}$ $512^{2}$ $256^{2}$ $256^{2}$
Format PNG PNG JPEG PNG, JPEG PNG, JPEG
Data: Our experiments are based on a set of $1,000$ real face images from the FFHQ dataset [19] dataset and another $2,000$ images created with generative AI models from the $DF^{3}$ dataset [17]. All images contain a single human face. Two AI generative models are considered, namely StyleGAN2 [20] and Latent Diffusion [29]. We also adopt two evaluation protocols from the $DF^{3}$ dataset [17]. This includes assessing the basic detection performance of raw data and evaluating the robustness of post-processed DeepFake data through mixed operations such as JPEG Compression, Gaussian Blur, face blending, adversarial attacks, and multi-image compression. Detailed information on the data used is given in Table 1. A few examples of the real and AI-generated images are shown in Fig. 3.
@@
Figure 4: Examples of GPT4V for DeepFake face detection. Left: Results for AI-generated images from the $DF^{3}$ dataset [17]. Right: Results for real faces from the FFHQ dataset [19]. The responses for AI-generated faces are labeled in pink, while those for the real faces are labeled in green. Both success (w/ Refer to caption) and failure (w/ Refer to caption) cases are shown.
@@
Performance Metrics: For each text-image prompt, we query the LLM multiple times and calculate a numerical score by averaging the results (No $=0$ and Yes $=1$). This approach offers two benefits. Firstly, it diminishes the variability in LLM responses to identical queries, attributable to the probabilistic nature of the underlying LLMs. Secondly, using numerical decision scores enables the application of performance metrics beyond mere accuracy, such as the area under the ROC (AUC) score. Compared to classification accuracies, the AUC score is less affected by imbalanced data, provides a more comprehensive performance evaluation, and allows us to compare the LLM’s performance with existing programmed detection methods. AUC score is a real number in $[0,1]$, with higher values corresponding to better performance. As the LLM may decline to respond to a query, another important performance metric is the rejection rate, which measures the fraction of queries that the LLM declines. We also report the single-class accuracy at the fixed threshold of 0.5.
@@ 4.1 Qualitative and Quantitative Results
The quantitative results corroborate this observation. Fig. 6 shows the receiver operational curves (ROCs) and the corresponding AUC scores obtained using API calls (as described in Section 3) over the evaluation dataset with the same prompt – GPT4V has a $79.5\%$ AUC on raw latent diffusion-generated face images and $77.2\%$ AUC score on StyleGAN-generated face images. The performance confirms that the GPT4V model obviously did not make random guesses on this task (corresponding to a ROC as a diagonal line and a $50\%$ AUC score). Compared to the GPT4V model, Gemini shows a slight decrease in performance.
@@
Figure 6: ROC curves of GPT4V and Gemini 1.0 Pro on the DeepFake detection based on averaging the predictions of five rounds of queries, (a) on raw data, (b) on post-processed Deepfake data.
@@
Table 3: Comparison of single-class Accuracy (%) in detecting DeepFake faces. ‘SG2’ stands for the StyleGAN2 model, and ‘LD’ represents the Latent Diffusion model.
@@
On the other hand, we note that most errors of GPT4V occur on real images – per Table 3, the classification accuracies on real images are around $50\%$, drastically different from those of AI-generated images, which is above $90\%$. Some intuitions can be obtained when we examine the real face images for these error cases, as shown in Fig. 4. These cases include semantic features unusual for “typical” face images. For instance, different age group (baby in the first case) or unique hair color (second case) to style (third case). This suggests that the semantic abnormality identified by GPT4V may not be specific to DeepFake faces. This problem may be solved by refining the model. In contrast, the Gemini model achieves a classification accuracy of $83.3\%$ on real images, dropping to around $50\%$ on generated faces. The examples in Fig. 5 show that the Gemini model’s response lacks rationality in analyzing the synthesis artifacts.
@@ 4.3 Improvements
So far, we have only tested simple queries. It has been demonstrated that using better prompts constructed with chain-of-thought prompts [36], few-shot prompting [41], which provide step-to-step guides in an interactive conversation with the LLM can elicit more relevant responses. The API interface of GPT4V and Gemini 1.0 does not allow multiple rounds of dialog because no consistent context is stored across API calls. Therefore, such interactive guidance can only be used with the web interface through a manual interaction, i.e., they cannot be automated using API calls. We provide two exploratory approaches in Fig. 10: one employs decomposed local images, a method we refer to as decomposition-based prompting, and the other utilizes a few-shot prompting technique. By supplying decomposed parts of images, we direct the model’s attention towards finer local patterns and reveal subtle visual anomalies. Further, using few-shot prompting with brief synthesis instructions imparts crucial forensic knowledge to the model, enabling the model to correctly classify two samples that were previously misidentified. These initial results suggest that using more crafted prompts can improve performance. However, we will wait for the LLMs to enable consistent API calls.
@@ 5 Conclusion
In this study, we investigate the potential of leveraging multimodal LLMs for tasks related to media forensics. Our research will broaden the application of multimodal LLMs to include a wider array of media forms, particularly focusing on video analysis. Rather than simply applying image-based detection techniques to video frames, a more integrated approach would involve direct video content processing. Additionally, we aim to enhance the detection of text-image mis-contextualization [15], where images and text are misleadingly paired to spread false information. Future endeavors will also explore developing more sophisticated prompting strategies and integrating these models with conventional signal or data-driven detection techniques.
@@
Acknowledgement. Siwei Lyu is supported by U.S. National Science Foundation (NSF) under grant SaTC-2153112. Xin Li is supported by U.S. NSF under grant CCSS-2348046 and SUNY-Albany start-up funds. Baoyuan Wu is supported by National Natural Science Foundation of China under grant No.62076213, Shenzhen Science and Technology Program under grant No.RCYX20210609103057050, and the Longgang District Key Laboratory of Intelligent Digital Economy Security.